What is the range of the function f(x)=4x-3 if the domain is {-1.0.2}?
Answer by stanbon(75887)

What is the range of the function f(x)=4x-3 if the domain is {-1.0.2}?
----
f(-1) = -4-3 = -7
f(0) = 0-3 = -3
f(2) = 8-3 = 5
-----
Range = {-7,-3,5}
======================
Cheers,
Stan H.
======